an angry construcion worker throws his wrench downward from a height of 128 feet with an intital velocity of 32 feet per second. the height of the wrench above the ground after t seconds is given by s(t)=-16t^2-32t+128.
A) what is the heigh of the wrench after 1 second ?
B)how long does it take for the wrench to reach the ground

Answer by nerdybill(7384)

s(t)=-16t^2-32t+128
.
a) simply plug substitute 1 for a t and solve:
s(t)=-16t^2-32t+128
s(1)=-16(1)^2-32(1)+128
s(1)=-16-32+128
s(1)=-48+128
s(1)= 80 feet
.
b) substitute 0 for s(t) and solve for t:
s(t)=-16t^2-32t+128
0=-16t^2-32t+128
Divide both sides by -16:
0=t^2+2t-8
Factor:
0=(t+4)(t-2)
.
t = {-4,2}
.
Toss out the negative solution leaves with:
time for wrench to reach ground = 2 seconds